OpenStudy (anonymous):

Steve travels 3 times faster than Mike. Traveling in opposite directions, they are 170 miles apart after 2.5 hours. Find their rates of travel.

OpenStudy (callisto):

let x the the rate of travel of Mike, (3x+x)*2.5 = 170 4x=68 x=17miles/hour therefore Mike travels at 17miles/hour while Steve travels at 51miles/hour

OpenStudy (anonymous):

How did you figure it out? Every time I've done it, the computer says it's wrong. I think I'm missing a step.

OpenStudy (callisto):

what do you mean?

OpenStudy (callisto):

(3x+x)*2.5 = 170 is actually from speed x time = distance and they travel in opposite direction so i just take the time as the common factor

OpenStudy (anonymous):

let speed of Mike = x mph Then speed of Steve = 3x mph In 2.5 hours Mike covers distance = x * 2.5 = 2.5x Steve covers distance = 3x * 2.5 = 7.5x ATQ 2.5x + 7.5x = 170 10x = 170 so x = 17 mph so Mike travels at 17 mph and Steve travels at 3*17 = 51 mph

OpenStudy (anonymous):

Okay, I get it now. Thank you.

OpenStudy (anonymous):

U r welcome....☺